stocks c and m each have an expected return of 9%, a beta of .90, and standard deviation of 30%. the returns on the two stocks have a correlation of .50. portfolio p has 60% in stock c and 40% in m. which is true?

a) p standard dev less than 30%

b) p standard dev equal to 30%

c) p beta is less than .90

d) p beta is greater than .90

e) p expected return greater than 9%
Answer #1

Beta of P will be 0.9 itself as both C and M have Beta of 0.9.

SD of portfolio = (W1^2 * SD1^2 + W2^2 * SD2^2 + 2* W1* W2 * SD1 * SD2* Correl(C,M))^1/2

=(0.6^2 * 0.3^2 + 0.4^2*0.3^2 + 2* 0.6*0.4 *0.3*0.3* 0.5)^1/2 = 26.15%

So p standard deviation is less than 30%.

Answer is a) p standard dev less than 30%